# Visual-Based Forklift Learning System Enabling Zero-Shot Sim2Real Without Real-World Data

### v1

@@ IV-D DRL on Isaac Sim
We used “OmniIsaacGymEnvs,” a parallel learning framework that is based on “rl_games” and available in Isaac Sim [33, 34]. This tool allows for constructing multiple environments on the same ground plane, improving the sampling efficiency. The randomization of the colors for pallets and other objects was handled independently for each environment for domain randomization. We synchronized task resets and shared the values for lighting and floor color randomization across all environments to simplify the implementation.
@@ IV-E Decision policy for loading
To perform loading, it is crucial to verify that the approach to the pallet has been successful. In this study, we designed a decision policy that activates when the forklift is stationary. The decision policy determines whether to lift the forks after the forklift approaches the pallet using the approach policy obtained through DRL. This policy was trained using binary classification through supervised learning. The training dataset was generated by collecting camera images from successful and unsuccessful attempts within the Isaac Sim environment, as shown in Fig. 7. Consequently, no real-world data were used to create this dataset.
@@ V Real demonstration
This section presents the results of deploying our method in a real environment. As our method employs zero-shot sim2real, the policies trained in the digital environment was directly transferred to the real environment without any modifications.
@@ V-A Setup
We trained the policies in the digital environment described in Section IV using the method outlined in Section III. The ground control computer with 96 GB of RAM, an Intel® Xeon Gold 6146 3.2 GHz processor, and an NVIDIA RTX A6000 was used. The task was performed by positioning the forklift within a triangular area, similar to that during the training in the digital environment, as shown in Fig. 4. In this demonstration, we repeatedly positioned the forklift at an initial position, directed it to approach the pallet, and manually returned it to a new starting position after each task.
The decision policy for determining whether to lift the forks was triggered 3s after the forklift came to a stop. In the case of a successful decision of approach, the forklift followed a sequence of actions: lifting the forks, reversing, and lowering the forks to unload the pallet. Successful completion of the task was dependent on executing the sequence without dropping the pallet load.
We compared the results of our method with those of human-operated trials to demonstrate the difficulty of the task. Human operators controlled the forklift using a joystick, relying on the same camera images and velocity information as the approach policy. Before the experiment, the operators were provided with a brief explanation and a 3-min practice session. As hydraulic operation requires experience, the experiment was considered successful if the operator could insert the forks at least two-thirds into the pallet.
@@ V-B Results
We performed the task 10 times with our learned policies and the three human operators. The success rates and average times of the pallet approach required for success are shown in Table II. The average success rate for human operators was 60% and the completion time was 16.5 s. Although Man C achieved a success rate of 90%, he employed a cautious driving technique, repeatedly stopping and checking the camera images frequently. This indicates that the task was not straightforward. Our method achieved a 60% success rate on the pallet approach task at the highest speed. Figure 8 shows the position and velocity of the forklift forks with our method during the 10 trials based on the motion capture system. This result suggests that our approach policy learned efficient behavior, accelerating initially and decelerating near the pallet. Furthermore, the success rate of loading decisions was 90%. This decision error occurred once when the forks were inserted only on one side. Snapshots of a successful attempt are presented in Fig. 9.
@@ VI Conclusions and future work
We have proposed a learning system for a counterbalance forklift to perform pallet loading tasks using only cameras and velocity data. In addition, we developed a 1/14-scale forklift with the same configuration as the 1/1-scale one. The forklift allows us to validate the trained policies for safety. Moreover, our method eliminates reliance on real-world data by utilizing a photorealistic environment and domain randomization. Our approach policy achieved a 60% success rate through efficient behavior in loading tasks in real-world experiments. Additionally, using a decision policy learned from a dataset constructed with a digital environment, the forklift could perform loading tasks based on the success or failure of its approach to the pallet. These policies were implemented through zero-shot sim2real, requiring no heuristic adjustments. Based on these results, this study presented a practical method for applying learning-based control to counterbalance forklifts, offering significant potential to advance automation across various industries.
Future research will focus on improving the accuracy using the results of various existing studies and validating the findings on a full-scale forklift. In addition, as a higher-level task, we will explore integrating control systems and task allocation strategies to automate forklift tasks fully.

### v2

@@ IV-D DRL on Isaac Sim
We used “OmniIsaacGymEnvs,” which is a parallel learning framework that is based on “rl_games” and is available in Isaac Sim [33, 34]. This tool enables the construction of multiple environments on the same ground plane, thereby improving the sampling efficiency. The randomization of the colors for the pallets and other objects was handled independently for each environment for domain randomization. We synchronized the task resets and shared the values for lighting and floor color randomization across all environments to simplify the implementation.
@@ IV-E Decision policy for loading
To perform loading, verifying that the approach to the pallet has been successful is crucial. In this study, we designed a decision policy that activates when the forklift is stationary. The decision policy determines whether to lift the forks after the forklift approaches the pallet using the approach policy obtained through DRL. This policy was trained using binary classification through supervised learning. The training dataset was generated by collecting camera images from successful and unsuccessful attempts within the Isaac Sim environment, as shown in Fig. 7. Consequently, no real-world data were used to create this dataset.
@@ V Real demonstration
This section presents the results of deploying our method in a real environment. As our method employs zero-shot sim2real, the policies trained in the digital environment were transferred directly to the real environment.
@@ V-A Setup
We trained the policies in the digital environment described in Section IV using the method outlined in Section III. A ground control computer with 96 GB of RAM, an Intel® Xeon Gold 6146 3.2 GHz processor, and an NVIDIA RTX A6000 was used. The task was performed by positioning the forklift within a triangular area, similar to that during the training in the digital environment, as shown in Fig. 4. In this demonstration, we repeatedly positioned the forklift at an initial position, directed it to approach the pallet, and manually returned it to a new starting position after each task.
The decision policy for determining whether to lift the forks was triggered 3 s after the forklift came to a stop. In the case of a successful approach decision, the forklift followed a sequence of actions: lifting the forks, reversing, and lowering the forks to unload the pallet. Successful completion of the task was dependent on executing the sequence without dropping the pallet load.
We compared the results of our method with those of human-operated trials to demonstrate the difficulty of the task. Human operators controlled the forklift using a joystick, relying on the same camera images and velocity information as the approach policy. The operators were provided with a brief explanation and a 3-min practice session before the experiment. As hydraulic operation requires experience, the experiment was considered successful if the operator could insert the forks at least two-thirds into the pallet.
@@ V-B Results
We performed the task 10 times with our learned policies and the three human operators. The success rates and average times of the pallet approach required for success are shown in Table II. The average success rate for human operators was 60% and the completion time was 16.5 s. Although Man C achieved a success rate of 90%, he employed a cautious driving technique, stopping repeatedly and checking the camera images frequently. This indicates that the task was not straightforward. Our method achieved a 60% success rate in the pallet approach task at the highest speed. Figure 8 shows the position and velocity of the forklift forks with our method during the 10 trials based on the motion capture system. This result suggests that our approach policy learned efficient behavior, accelerating initially and decelerating near the pallet. Furthermore, the success rate of the loading decisions was 90%. A decision error occurred once when the forks were inserted only on one side. Snapshots of a successful attempt are presented in Fig. 9.
@@ VI Conclusions and future work
We have proposed a learning system for a counterbalance forklift to perform pallet loading tasks using only cameras and velocity data. In addition, we developed a 1/14-scale forklift with the same configuration as the 1/1-scale one. The forklift allows us to validate the trained policies for safety. Moreover, our method eliminates reliance on real-world data by utilizing a photorealistic environment and domain randomization. Our approach policy achieved a 60% success rate through efficient behavior in loading tasks in real-world experiments. In addition, the forklift could perform loading tasks based on the success or failure of its approach to the pallet using a decision policy learned from a dataset constructed with a digital environment. These policies were implemented through zero-shot sim2real, requiring no heuristic adjustments. Based on these results, this study presented a practical method for applying learning-based control to counterbalance forklifts, offering significant potential to advance automation across various industries.
Future work will extend our training approach and environment to accommodate more diverse scenarios. In addition, we will incorporate insights from existing studies to improve the accuracy further and validate these results on a full-scale forklift.